What is the momentum of a 0.04kg bullet moving north at a rate of 1500m/s?



What are the given:




What is asked in the problem:



Formula to use.



Solution:

1
Expert's answer
2022-06-14T13:39:05-0400

Given:

"m=0.04\:\rm kg"

"v=1500\:\rm m/s"

Unknown:

"p"



"p=mv\\=0.04*1500=60\:\rm kg\cdot m/s"